HUBBLE SPACE TELESCOPE DAILY REPORT #1575 PERIOD COVERED: 0000Z (UTC) 02/21/96 - 0000Z (UTC) 02/22/96 Daily Status Report as of 053/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 6253 (Pure Parallel Archive Program) The WF/PC-2 was used to execute a pure parallel program taking = ultraviolet observations running concurrently with prime science = observations. The intent is to search for extra-galactic sources with = flux excesses at 3000 A0 and galactic sources with excesses at 1600 A0. = The observations were completed as planned, and no problems were reported. 1.2 Completed GHRS/1 5877 (The 11B to 10B Ratio in the Interstellar = Medium.) The High Resolution Spectrograph (Side One) was used to make = spectrographic observations of the bright star HD38771 in order to study = the foreground interstellar medium. This is to measure the abundance = ratio of two boron isotopes. The observations were completed as planned, = and no problems were reported. 1.3 Completed Two Sets of WF/PC-2 6187 (WF/PC-2 Cycle 5 Calibration: = Earth Flat) The WF/PC-2 was used to make Earth streak flat calibration images. = These streak flats will be used to produce flat fields, to map the OTA = illumination pattern and, in combination with other data, to produce new = superflats. The observations were completed as planned, and no problems = were reported. 1.4 Completed WF/PC-2 6344 (Low - Mass Binaries in the Hyades - = Completing the Survey) The WF/PC-2 was used to make observations of the low mass binary = star RHY369 (M-Class Dwarf Stars) which is located in the Hyades Cluster. = The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed WF/PC-2 6365 (A Snapshot Survey For Companions and = Clusters around Wolf-Rayet Stars) The WF/PC-2 was used to make snapshot observations of Wolf-Rayet = star WR125 in order to look for companion stars and clusters. Photometry = was performed on the Wolf-Rayet star and associated stars. The FGS = acquisition, at 053/073226Z, for this proposal failed. During the = acquisition attempt low PMT counts were observed. The HST was also in the = SAA during the period. Five observations were affected. HSTAR 5509 = documents these events. Otherwise, the observations completed as planned, = and no further anomalies were reported. 1.6 Completed FOS/RD 5852 (Lyman-alpha Forest Cloud Sizes and = Clustering at Z~0.5) The Faint Object Spectrograph (Red Side) was used to make = spectrographic observations of QSO MC1146+111 to study its Lyman-Alpha = forest and its relation to other nearby QSOs. The observations were = executed as scheduled, and no anomalies were noted. 1.7 Completed FOS/RD 6025 (Consequences of the Shoemaker-Levy Impact = With Jupiter) The Faint Object Spectrograph (Red Side) was used to make = observations of the planet Jupiter in order to study the consequences of = the impact of the comet Shoemaker-Leve with the planet. The observations = were completed as scheduled and no anomalies were noted. 1.8 Completed GHRS/2 5885 (GHRS- Observations of Candidate = Beta-Pictoris Circumstellar Gaseous Disks) The High Resolution Spectrograph (Side Two) was used to make = observation of star HD158643 which is a candidate to have Beta = Pictoris-like gaseous disks. The objective is to confirm the existence of = circumstellar gas, and to understand its origin. The observations were = completed as planned, and no anomalies were reported. 1.9 Completed WF/PC-2 6118 (HST Observations of Two Ultraluminous = Extragalactic Supernova Remnants) The WF/PC-2 was used to make observations of the ultraluminous = extragalactic supernova remnant NGC-4449. This is one of the brightest = supernova remnants known. Its high luminosity may be caused by = interaction with its local environment. The observations were completed = as scheduled, and no problems were reported. 1.10 Completed WF/PC-2 5971 (HST Medium Deep Survey) The WF/PC-2 was used to make a long exposure medium deep sky survey = observations at latitudes larger than 10 degrees taken in parallel with = the prime science observations. The observations were executed as = scheduled, and no anomalies were noted. 1.11 Completed FOS/BL 6166 (FOS Cycle 5: Location of Spectra and = Wavelength Calibration) The Faint Object Spectrograph (Blue Side) was used to make = wavelength calibration images to measure the stability of the FOS Blue = Side wavelength scale and spectral locations. A Taled calibration = observation was also taken. The observations were completed as planned, = and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 9 Successful : 8 One acquisition, at 053/073226Z, failed. Low PMT counts were = observed, and the acquisition did occur in the SAA. HSTAR 5509 was = written. Scheduled Reacquisitions : 5 Successful : 5 2.2 FHST Updates: Scheduled : 21 Successful : 21 2.3 Operations Notes: ESTR #2 remains in a powered off state. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH